DIRC Operations Weekly Report

Period: Saturday 2006/06/17 - Friday 2006/06/23

Oncall: Rolf Andreassen, Jerome Bequilleux
Fast Monitoring: Rolf Andreassen, Jerome Bequilleux

File location: $BFROOT/www/Detector/DIRC/WeeklyReports/report_20060623.html
Plot directory: $BFROOT/www/Detector/DIRC/WeeklyReports/Plots/report_20060623 (empty)

Contents


BaBar-PEP news

PEP peak lumi: 10.2*1033 (previous week: 10.2*1033)

BaBar integrated lumi: 2.831 fb-1 (previous week: 4.049 fb-1)

Run 5b PEP delivery: 62.49 fb-1

Run 5 total delivery: 120.72 fb-1 Run 5 total BaBar recorded: 116.35 fb-1 including 9.82 fb-1 off-peak data
Note : No PEP-BaBar meeting this week; numbers in red are estimated 'by hand' from plots. Numbers in black are official.

PEP activities this week:

  • Tuesday 20th : Beam tests to track down horizontal instability in LER; most of the day shift.
  • Thursday 22nd : Problems have been traced to failed components in absorptive feedback filters for the LER. Work to replace them, several hours down.
  • Friday 23rd : A circuit breaker was broken during repairs to a water pump. Shutdown for some hours.

    PEP activities planned next week :

  • Foreseen work opportunities (PEP down periods):

  • About two hours starting at 1100 on Monday, while the UPS is being repaired. Requires Matt to disable and enable SOB dump valve. At the same time PEP will replace some LER filters that are sickening.

    BaBar activities/issues this week:

  • DIRC activities

    See DIRC logbook for details
  • Nicolas adjusted the scaler-alarm histograms in the daily background monitoring, to take into account the new alarm values. The difference is very clear, percentage of shift in yellow alarm goes from 10 (Swing 19th) to 2 (Owl 20th).
  • Wed 21 / Thu 22 : A discrepancy was discovered in the temperatures reported by epics and the Honeywell box for the water temperature. Jerry suggested that this is due to offset drift, and gave a new formula for the analog-digital conversion, to compensate.
  • The flow in the DIRC water chiller has been around 4.4 all week; nominal is 4.8 +- 0.7. Water level seems normal; not clear what is causing this. Will keep an eye on it.

    Downtime due to the DIRC: None.

    Backgrounds

    Background summary plot

    Scheduled activities

    Pending Work

    Links